### v1

# Client-wise Modality Selection for Balanced Multi-modal Federated Learning

Abstract
Selecting proper clients to participate in the iterative federated learning (FL) rounds is critical to effectively harness a broad range of distributed datasets. Existing client selection methods simply consider the variability among FL clients with uni-modal data, however, have yet to consider clients with multi-modalities. We reveal that traditional client selection scheme in MFL may suffer from a severe modality-level bias, which impedes the collaborative exploitation of multi-modal data, leading to insufficient local data exploration and global aggregation. To tackle this challenge, we propose a Client-wise Modality Selection scheme for MFL (CMSFed) that can comprehensively utilize information from each modality via avoiding such client selection bias caused by modality imbalance. Specifically, in each MFL round, the local data from different modalities are selectively employed to participate in local training and aggregation to mitigate potential modality imbalance of the global model. To approximate the fully aggregated model update in a balanced way, we introduce a novel local training loss function to enhance the weak modality and align the divergent feature spaces caused by inconsistent modality adoption strategies for different clients simultaneously. Then, a modality-level gradient decoupling method is designed to derive respective submodular functions to maintain the gradient diversity during the selection progress and balance MFL according to local modality imbalance in each iteration. Our extensive experiments showcase the superiority of CMSFed over baselines and its effectiveness in multi-modal data exploitation.
1 Introduction
Figure 1: Traditional client selection in FL aims to sample a client subset in each round. The selected clients use all their local data for training and aggregation. To exploit different modalities comprehensively, we propose the client-wise modality selection for MFL, which considers the local data from each modality as the sampling unit rather than the client. Only the selected modality data from a client would join in training and aggregation.
Federated learning (FL) [20] aims to collaboratively learn data that has been collected by, and resides on, a number of remote devices or servers. FL stands to develop top-performing models by aggregating knowledge from numerous edge clients [6], which relies on the iterative interaction among participating clients and the server. However, comprehensively employing the information from all clients can be exceptionally difficult due to the client heterogeneity and resource limitations [8, 27].
Random sampling [32, 19] from available clients is widely used in FL to satisfy some practical restrictions, e.g., limited communication bandwidth [23, 31] and computing capacities [13]. To improve the information exploitation of all clients, extensive research has been conducted on effective client selection strategies. One viable solution [24] is to gather as much local data as possible in each round while in contrast, some alternative studies [1] focus on collecting representative information by sampling a subset of clients. However, previous methods have only considered the knowledge discrepancies across different clients, while demonstrating notable performance in uni-modal FL, yet, their effectiveness diminishes when dealing with clients with multi-modal data as the inter-modality interactions during the FL training is neglected. The investigation into fine-grained and sophisticated data selection mechanism for multi-modal FL (MFL) is currently under-explored.
In this paper, we reveal that traditional client selection approaches may suffer from a severe modality-level bias in MFL. Due to the inconsistent learning progress of different modalities, i.e., the modality imbalance, client-level selection bias tends to favor information from a dominant modality while deprioritize the weak modality, which impedes the collaborative exploitation of distributed multi-modal data, leading to insufficient local data exploration and global aggregation.
To address such issue in MFL, we propose a novel Client-wise Modality Selection scheme for MFL (CMSFed), which chooses the appropriate modalities combination from clients in each round to alleviate the modality-level bias and realize full exploitation of each modality via considering the interactions between different modalities as illustrated in Fig. 1. Only part of, not all, modalities, will be selected for some clients compared with traditional client selection. Specifically, in each communication round, we intend to gather a data subset by selectively employing different modalities from each client, the aggregated model update stemmed from that approximates the fully aggregated update which fulfills the potential of each modality. Inspired from [1] for diverse client selection, we build the modality selection formula, namely the min-form of the facility location function [5] evaluated on the collected data subset, to seek for the representative information for the global model. To further release the ability of each modality and reduce the training cost, we propose a modality-level gradient decoupling with a modal-balanced local training loss to build two separated submodular functions and then modulate their optimization process for unbiased modality selection. As a result, inconsistent modality adoption strategies may lead to different feature spaces for different clients, which exacerbates the difficulties in accomplishing an accurate and generalized aggregated model. Therefore, the global prototype is introduced for the weak modality to alleviate the heterogeneous feature spaces. To summarize, our contributions in this paper are as follows:
  • We scrutinize the client selection scheme in MFL and reveal that naive sampling clients may introduce a strong modality-level bias which causes insufficient utilization of multi-modal data.
  • We propose a novel client-wise modality selection scheme for MFL (CMSFed) to comprehensively exploit all modalities via a enhancing loss and gradient decoupling to balance MFL.
  • We conduct comprehensive experiments and demonstrate that 1) CMSFed can achieve considerable improvements over baselines; 2) CMSFed indeed achieves a more balanced exploitation of distributed multimodal data.
2 Background and Related Works
(a) Multi-modal
(b) Uni-audio
(c) Uni-visual
Figure 2: Performance of client selection strategies on CREMA-D with uniformly distributed data on 20 clients. FedAvg randomly selects clients while DivFL selects clients with diversity. ‘FedAvg-0.8’ means 80% of clients have data in both audio and visual modalities, while 20% of clients only have data in random one modality. ‘-0.5’ and ‘-0.2’ follow the similar definition. (a) Performance of global multi-modal model. (b) Performance of global uni-audio model. (c) Performance of global uni-visual model.
2.1 Multi-modal federated learning
In MFL, each client may contain one or more modalities of data. Without loss of generality, we consider two input modalities, referred to as $A$ and $I$ in MFL. There are a set $V$ of $N$ clients, $V=\left[N\right]$, respectively owning datasets $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},y\right\}$ composed of $n_{i}$ pair samples, $i\in\left[N\right]$. A typical federated learning objective is the average of each client’s local loss function:
$\vspace{-8pt}\min f\left(\theta\right)=\sum_{k=1}^{N}{p_{k}F_{k}\left(\theta\right)}$ (1)
where $\theta=\left\{\theta^{A},\theta^{I},\omega\right\}$ denotes the model parameters. $\theta^{A}$ and $\theta^{I}$ represent the parameters of encoders of modality $A$ and $I$. $\omega$ is the parameters of fusion classifier. $p_{k}$ is a pre-defined weight. $F_{k}$ is each client’s local loss.
Statistical heterogeneity [14, 17] is a widely concerned challenge in uni-modal FL. To tackle this issue, FedProx [18] uses a proximal term to stabilize model aggregation. FedProto [29] shares class prototypes to regularize the learning of local models. In MFL, modality incongruity [38, 37] (clients consist of different modalities combinations), as well as statistical heterogeneity [35], are all considered. Yu et al. [37] propose CreamFL to align the representations between different clients and different modalities via communicating knowledge on a public dataset. Chen et al. [3] introduce FedMSplit to split local models into several components and aggregate them by their correlations. However, they still focus on the heterogeneity, but ignore the interaction between private data of different modalities, which limits their information exploitation.
2.2 Client selection and submodular function
Client selection [7, 36] is a critical issue for FL especially when the communication cost with all devices is prohibitively high, which has been extensively studied in uni-modal FL. Cho et al. [4] propose Power-of-Choice to select clients with largets local loss. Balakrishnan et al. [1] propose to select a subset of clients with great diversity.
Diverse client selection via submodularity. Maximizing a submodular function is reported to improve the diversity and reduce the redundancy of a subset. This property makes it appropriate for client selection in FL. If a function $F$ is submodular, it should satisfy: given a finite ground set $V$ of size $N$, $F\left(A\cup\left\{v\right\}\right)-F\left(A\right)\geqslant F\left(B\cup\left\{v\right\}\right)-F\left(B\right)$, for any $A\subseteq B\subseteq V$ and $v\in V\backslash B$. The marginal utility of an element $v$ w.r.t. a subset $A$ is denoted as $F\left(v|A\right)=F\left(A\cup\left\{v\right\}\right)-F\left(A\right)$, which can represent the importance of $v$ to $A$. The client selection via submodular maximization can be expressed following [1]: find a subset $S$ of clients whose aggregated gradients can approximate the full aggregation from all clients:
$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}$ $=\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v^{\sigma\left(k\right)}\right)\right]}$ (2)
$+\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}$
where $\sigma$ maps $V\rightarrow S$ and the gradient $\nabla F_{k}\left(v^{k}\right)$ from client $k$ is approximated by the gradient from a selected client $\sigma\left(k\right)\in S$. For $i\in S$, let $C_{i}\triangleq\left\{k\in V|\sigma\left(k\right)=i\right\}$, and therefore $\gamma_{i}\triangleq|C_{i}|$. Take the norms and apply triangular inequality after subtracting the second term from both sides, we can obtain a relaxed objective $G\left(S\right)$ for minimizing the approximation error:
$\begin{split}&\left\|\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}-\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}\right\|\leqslant\\ &\sum_{k\in\left[N\right]}{\underset{i\in S}{\min}\left\|\nabla F_{k}\left(v^{k}\right)-\nabla F_{i}\left(v^{i}\right)\right\|\triangleq G\left(S\right)}\end{split}$ (3)
Minimizing $G\left(S\right)$ can be seen as maximizing the well-known submodular function, i.e., the facility location function [5]. The submodular maximizing problem is NP-hard but can be approximated via the greedy [22] or stochastic greedy algorithm [21]:
$S\leftarrow S\cup k^{*},$ (4)
$k^{*}\in$ $\underset{k\in\mathrm{rand}\left(V\backslash S,\mathrm{size}=s\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S\right)-\bar{G}\left(\left\{k\right\}\cup S\right)\right]$
$\bar{G}$ represents a constant minus the negation of $G$.
Although these methods make great improvement in uni-modal FL, the selection strategy is under-explored in MFL and we reveal that traditional client selection approaches fall into a severe bias in MFL.
2.3 Imbalanced multi-modal learning
Modality imbalance indicates the inconsistent learning progress of different modalities in multi-modal learning [33, 12]. Peng et al. [26] propose OGM-GE to alleviate the inhibitory effect on weak modality by slowing down the dominant modality. Fan et al. [9] further build a non-parametric classifier by class centroids to adjust the update direction of weak modality. In this paper, we aim to power each modality of all clients by a meticulously designed modality selection strategy in each round of training.
3 Observations and Analysis
In this section, we analyze the behavior of different client sampling strategies in MFL and the interactions between different modalities. As shown in Fig. 2, we compare a well-designed client selection method (DivFL) with random sampling (FedAvg) on a multi-modal dataset. The results in Fig. 1(a) show that the improvement from DivFL is severely limited although it is terrific in uni-modal FL. The results about how each modality behaves are illustrated in Figs. 1(b) and 1(c). It is apparent that audio modality performs much better than visual modality and the optimised client selection options from DivFL generally focus on audio modality (audio performance of DivFL has significant improvement over FedAvg while no noticeable difference in visual).
The performance gap between audio and visual can be attributed to ‘modality imbalance’ [33], which indicates the inconsistent learning paces in different modalities. Therefore, the imbalance nature (multi-modal learning is dominated by the modality who learns fast) [34] in MFL leads to a strong modality-level selection bias, i.e., client selection strategies tend to select clients that make the dominant modality aggregation better but neglect the weak modality
To further analyze the interactions between different modalities, we randomly discard the data of one modality (audio or visual) on clients with certain probabilities (FedAvg-0.8, -0.5, -0.2) as demonstrated in Fig. 1(a). It is clear that randomly discarding one modality data for part of clients does not necessarily lead to performance degradation of the global model, but may bring performance improvement. The reason behind can be found in Figs. 1(b) and 1(c), where the gap between audio and visual modalities gradually narrows as the discarding probability increases until there is no stark disparity (in FedAvg-0.2). This phenomenon is consistent with the findings in [26]: the dominant modality hinders the learning process of weak modality. Therefore, the ability of visual modality is released and absorbed by the global model through aggregation when the audio data in some clients is abandoned.
The above analyses suggest that traditional client selection methods are not applicable in MFL due to the modality imbalance nature, and the potential of each modality can be further explored by engaging appropriate modalities in training. Therefore, we intend to design a new modality selection scheme in MFL to address the modality imbalance and information underutilization from different modalities.
4 Method
In this section, we introduce CMSFed, a diverse and balanced modality selection strategy to realize more efficient data utilization. We first derive a combinatorial objective for modality selection in MFL via the gradient approximation with the fully aggregated update. Then we propose a modality-level gradient decoupling with the enhancing gradient on weak modality to build two separated submodular functions and balance MFL.
4.1 Client-wise modality selection
Firstly, we rewrite Eq. 2 for MFL as:
$\begin{split}&\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)}=\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)}+\\ &\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v_{A}^{\sigma\left(k\right)},v_{I}^{\sigma\left(k\right)}\right)\right]}\end{split}$ (5)
where $v_{A}^{k}$ and $v_{I}^{k}$ indicate the corresponding modal data in client $k$.
The above equation is still the client selection formulation, while we could build the modality selection formula based on it as (data superscripts are omitted for simplicity):
$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}$ (6)
$=\sum_{k\in\left[N\right]}{\left[\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\nabla F_{\sigma_{M}\left(k\right)}\left(v_{A},v_{I}\right)\\ -\nabla F_{\sigma_{A}\left(k\right)}\left(v_{A}\right)-\nabla F_{\sigma_{I}\left(k\right)}\left(v_{I}\right)\\ \end{array}\right]}$
$+\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}+\sum_{k\in S_{A}}{\gamma_{k}^{A}\nabla F_{k}\left(v_{A}\right)}$
$+\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}$
where $\sigma_{M}$, $\sigma_{A}$ and $\sigma_{I}$ map $V\rightarrow S_{M},S_{A},S_{I}$, the client sets who use multi-modal, $A$ modal, and $I$ modal data for training respectively, and $S_{M}\cap S_{A}=S_{A}\cap S_{I}=S_{M}\cap S_{I}=\oslash$. The full gradient aggregation is represented by the customized update from multi-modal and uni-modal clients. Assume modality $I$ is weak here, the relaxed objective for modality selection should be:
$\left\|\begin{array}[]{c}\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}-\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}\\ -\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}\\ \end{array}\right\|$ (7)
$\leqslant\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\\ \nabla F_{i}\left(v_{A},v_{I}\right)-\nabla F_{j}\left(v_{I}\right)\\ \end{array}\right\|}$
uni-$A$ modal is omitted here as the multi-modal gradient is dominated by $A$, which means we do not need to select uni-$A$ clients for its acceleration. However, minimizing this objective, which is similar to client selection in Eq. 3, cannot leverage sufficient information from each modality since the approximation objective (the left-hand side of Eq. 6) still suffers from serious modality imbalance. How to finding a new full gradient aggregation that can exploit all modalities’ information is a great challenge. In addition to this, there also exist other challenges making Eq. 7 difficult to be effectively optimized in practice:
  • We need to build the relationship (a gradient similarity matrix) between fully aggregated gradient and local gradient for solving Eq. 7. In uni-modal FL, a client only needs to train once to get gradients while in MFL, each client should perform local update twice (with multi-modal and weak-modal data respectively) to determine whether to choose multi-modal or uni-modal data for training. This inevitably leads to almost double the local training cost.
  • Applying modality selection would cause different modality adoption strategies, namely modality heterogeneity, which results in divergent feature spaces across clients, especially on weak modality.
To address the above challenges, we propose a modality-level gradient decoupling method with a balanced local loss, which not only leads to comprehensive information exploitation but also notably saves training overhead.
4.2 Modality-level gradient decoupling
We focus on the classification tasks in this paper, where cross-entropy (CE) loss is widely used. Hence, the loss for multi-modal and uni-modal learning should be:
$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)$ (8)
$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)$
The naïve multi-modal loss falls into modality imbalance. Although obtaining a totally balanced solution seems impractical for multi-modal learning, we can find a surrogate for it: here we enhance the weak modality by equipping the naïve losses with the prototypical cross-entropy (PCE) loss $\beta^{k}\mathcal{L}_{P}^{k}$ [9] to correct gradient directions of weak modality:
$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:\tilde{F}_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)$ (9)
$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:\tilde{F}_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)+\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)$
where $\mathcal{L}_{P}^{k}$ is calculated based on the local prototypes $c^{I}$:
$\mathcal{L}_{P}^{k}\left(v_{I}\right)=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{I}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{I}\right)\right)}}\right]$ (10)
where $d\left(\cdot,\cdot\right)$ is the distance function (Euclidean distance), $z_{i}^{I}$ is the feature of $x_{i}^{I}$. $Y$ is the class number. The calculation details for $c^{I}$, $\beta^{k}$ are in Appendix. Applying $\mathcal{L}_{P}^{k}$ on uni-modal loss is also essential, which can improve this modality by clustering features tightly.
Bring Eq. 9 to Eq. 7 and we can get the new objective:
$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\tilde{F}_{k}\left(v_{A},v_{I}\right)-\\ \nabla\tilde{F}_{i}\left(v_{A},v_{I}\right)-\nabla\tilde{F}_{j}\left(v_{I}\right)\\ \end{array}\right\|}=$ (11)
$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\nabla\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{i}\left(v_{A},v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{P}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{P}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
$\leqslant\sum_{k\in\left[N\right]}{\underset{i\in S_{M}}{\min}\left\|\nabla\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)-\nabla\mathcal{L}_{CE}^{i}\left(v_{A},v_{I}\right)\right\|}$
$+\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{P}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{P}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
$\triangleq G\left(S_{M}\right)+G\left(S_{M}\cup S_{I}\right)$
Firstly, replacing Eq. 8 with Eq. 9 can enhance the weak modality on the local side. However, the right-hand side of the first equation is still a complex issue. Then, we decouple this objective into two submodular functions according to triangle inequality and the full gradient approximation is divided into two parts: the first part uses selected multi-modal CE gradient to fit fully multi-modal CE gradient aggregation and the second part approximates the fully multi-modal PCE gradient aggregation via selected uni-modal CE gradient and both selected multi- and uni-modal PCE gradient. Two gradient similarity matrices are required for two parts. The modality-level gradient decoupling not only reduces algorithm complexity (convert joint selection to two selection problems) but also highly lowers training cost as each selected client only needs to perform once local training according to its selected modality type to update gradient similarity matrices (both matrices for multi-modal clients, only the second matrix for uni-modal clients).
After modality selection, the gradient of $\theta^{I}$ from each client varies dramatically since there exists an obvious difference between the multi-modal learning gradient and uni-modal learning gradient for the weak modality, as discussed in [9]. To address this heterogeneity, we replace the local prototype in Eq. 10 with the global prototype $c^{GI}$:
$\mathcal{L}_{GP}^{k}\left(v_{I}\right)=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{GI}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{GI}\right)\right)}}\right]$ (12)
Applying the global PCE loss to Eq. 9 can enhance local weak modality and align the feature spaces of various clients simultaneously.
Then, we can perform modality selection with the stochastic greedy algorithm for two submodular functions:
$S_{M}\leftarrow S_{M}\cup k_{1}^{*},$ (13)
$k_{1}^{*}\in$ $\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S_{M}\right)-\bar{G}\left(\left\{k\right\}\cup S_{M}\right)\right]$
Since the second submodular function involves both $S_{M}$ and $S_{I}$, we add a simple yet effective conflict resolution strategy to ensure $S_{M}\cap S_{I}=\oslash$ as well as, more importantly, balance the learning of different modalities:
$\begin{cases}if\,\,k_{1}^{*}=k_{2}^{*},S_{M}\cup k_{2}^{*};\\ if\,\,k_{1}^{*}\neq k_{2}^{*},\left\{\begin{array}[]{c}S_{I}\cup k_{2}^{*},if\,\,\rho_{I}^{k}>\chi\\ S_{M}\cup k_{2}^{*},if\,\,\rho_{I}^{k}\leqslant\chi\\ \end{array}\right.\\ \end{cases}$ (14)
$k_{2}^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\begin{array}[]{c}\bar{G}\left(S_{M}\cup S_{I}\right)-\\ \bar{G}\left(\left\{k\right\}\cup S_{M}\cup S_{I}\right)\\ \end{array}\right]$
where $\rho_{I}^{k}$ is the imbalanced ratio of client $k$ as in [9], which is the quotient of average ground-truth logits from two modalities (See the Appendix for details). $\rho_{I}^{k}>1$ means audio is dominant while visual is dominant when $\rho_{I}^{k}<1$. $\chi$ is a hyper-parameter for threshold. It tends to select weak modality for severely imbalanced clients to alleviate imbalance, while selected balanced clients use multi-modal data to improve overall performance.
Input: Input data $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},y\right\}$, $i\in\left[N\right]$, initial model $\theta$, hyper-parameters $\chi$, global communication epochs $E$, $e=1$.
while $e<E$ do
       if $e=1$ then
             Send $\theta$ to all clients;
             Perform one-step local update for gradients, prototypes and $\rho_{I}^{k}$;
             Aggregate global prototypes and $\rho_{I}$;
      else
            Aggregate global model, prototypes and $\rho_{I}$;
      Select a subset of modalities $S_{M}$ and $S_{I}$ (or $S_{A}$) using Eqs. 13 and 14;
       Send $\theta$ to selected clients;
       foreach client in selected clients in parallel do
             Perform multi-modal learning in $S_{M}$ and uni-modal learning in $S_{I}$ (or $S_{A}$) by Eq. 9 with Eq. 12;
             Send gradients, prototypes and $\rho_{I}^{k}$ to server;
Algorithm 1 CMSFed.
(a) CREMA-D
(b) AVE
(c) CG-MNIST
Figure 3: Test accuracy of CMSFed compared with other baselines on CREMA-D, AVE and CG-MNIST under IID settings. CMSFed converges to more accurate solutions than all baselines in these three datasets.
(a) CREMA-D
(b) AVE
(c) CG-MNIST
Figure 4: Test accuracy of CMSFed compared with other baselines on CREMA-D, AVE and CG-MNIST under non-IID settings. CMSFed still makes remarkable improvement.
Discussion. (1) Mitigating the selection bias in our method are twofold: the PCE loss $\tilde{F}_{k}$ alleviates imbalance at local side and the selected uni-modal clients further promote balanced learning of global model. Meanwhile, the diversity coming from two submodular functions ensures the improvement for the global model. (2) The maximization of $\bar{G}\left(S_{M}\right)$ and $\bar{G}\left(S_{M}\cup S_{I}\right)$ can be viewed as the optimization for multi-modal learning and uni-weak-modal learning respectively. Imbalance ratio is adopted to keep their balance. (3) We assume $I$ is the weak modality above while in practice, we can determine the weak modality before modality selection via the aggregated global imbalance ratio $\rho_{I}$ with $\rho_{I}^{k}$ from all clients. Overall, the pseudo-code of CMSFed is provided in Algorithm 1. (4) Only the gradients, prototypes and $\rho_{I}^{k}$ participate in communication, so there is no privacy issue and similar communication overheads as in traditional FL.
5 Evaluation
5.1 Datasets and baselines
CREMA-D [2] is an audio-visual dataset for emotion recognition task, each video in which consists of both facial and acoustic emotional expressions. There are total 6 categories for emotional states: neutral, happy, sad, fear, disgust and anger. 7,442 clips in total are collected in this dataset. 6,698 samples are randomly chosen as the training set and the rest of 744 samples are the testing set.
AVE [30] is an audio-visual dataset for audio-visual event localization, which consists of 28 event classes and 4,143 10-second video clips with both auditory and visual tracks as well as second-level annotations. All the video clips are collected from YouTube. In our experiments, we aim to construct a labeled multi-modal classification dataset by extracting the frames from event-localized video segments and capturing the audio clips within the same segments. The training and validation splits of the dataset follow [30].
Colored-and-gray MNIST [15] is a synthetic dataset based on MNIST [16], and we denote it as CG-MNIST in this paper. Two kinds of images form a sample pair: a gray-scale image and a monochromatic image. In the training set, there are 60,000 sample pairs, and the monochromatic images are strongly color-correlated with their digit labels, In the validation set, the number of sample pairs is 10,000, while the monochromatic images are weakly color-correlated with their labels. The data synthesis method follows [34]. This dataset is used to prove the method’s effectiveness beyond audio-visual modality.
(a) Audio in CREMA-D
(b) Visual in CREMA-D
(c) Audio in AVE
(d) Visual in AVE
Figure 5: The performance of each modality compared with other baselines on CREMA-D and AVE under IID settings. CMSFed achieves comparable audio accuracy and significantly better visual performance simultaneously.
Baselines. We choose eight baselines for comparison from four categories: (1) three uni-modal FL methods designed for statistical heterogeneity are extended to multi-modal scenarios: FedAvg [20], FedProx [18] and FedProto [29]. (2) Integrating OGM-GE [26] and PMR [9], the solutions for modality imbalance, with FedAvg forms two MFL methods: FedOGM and FedPMR. (3) Two client selection method: Power-of-choice (pow-d) [4] and DivFL [1], evolved from its uni-modal version directly. (4) One MFL method, FedMSplit [3], especially designed for modality incongruity. Compared with these baselines, we demonstrate that an elaborate modality selection strategy is essential to realize comprehensive information exploitation in MFL.
5.2 Experimental settings
For CREMA-D and AVE, we use ResNet18 [10] as the backbone for both audio and visual modalities. Audio data is converted to a spectrogram of size 257x299 for CREMA-D and 257×1,004 for AVE. We randomly choose 3 frames and 4 frames to build image training sets for CREMA-D and AVE respectively. For CG-MNIST, we build a neural network with 4 convolution layers and 1 average pool layer as the encoder, following the setting as in [9]. We choose the simple yet effective fusion method, concatenation [25], to build fusion classifier for the three datasets. We set 20 clients for CREMA-D, AVE and 30 clients for CG-MNIST, while 5 clients are selected in each communication round for CREMA-D, AVE and 6 for CG-MNIST. For IID setting, training data is uniformly distributed to clients. For non-IID scenarios, we use Dirichlet distribution [11] $Dir\left(\alpha\right)$ to split data ($\alpha=3$ for CREMA-D and AVE, $\alpha=2$ for CG-MNIST). The optimizer is SGD [28] for all datasets. Learning rate is initialized at 1e-3 or 1e-2 for CEAMA-D and AVE or CG-MNIST and becomes 1e-4 or 1e-3 in the later training stage. The hyper-parameter $\chi$ is set to 1.5-2.5 according to datasets and settings. To complete stochastic greedy algorithm for Eqs. 13 and 14, we use the gradients from the selected clients at current round to update part of the similarity matrix, which is named “no-overheads” in [1]. Except for pow-d, DivFL and CMSFed, other baselines select clients randomly. Each client has two modal data by default if not specified.
5.3 Comparison with baselines
CMSFed effectively improves the performance. As shown in Fig. 3, we report the test accuracy versus the number of communication rounds for three datasets under IID data distributions. We observe that CMSFed can achieve significant improvement compared with all other baselines (by up to 5% on CREMA-D). CMSFed also realizes comparable or even faster convergence speeds in CREMA-D and CG-MNIST. The slow learning at the beginning in AVE is attributed to the need for more effort to explore the information of weak modality. Client sampling here (pow-d and DivFL) cannot fully exploit information for all modalities, making its improvement limited or even worse than FedAvg in CG-MNIST. Although FedOGM and FedPMR accomplish modest improvement because of their ability to alleviate modality imbalance, they are not as good as CMSFed since they do not consider the overall performance of aggregated model. Traditional uni-modal FL methods for statistical heterogeneity (e.g. FedProx) and MFL method for modality incongruity (FedMSplit) only obtain slight improvement. The results for non-IID data distributions are demonstrated in Fig. 4. The overall trends are similar to the results in Fig. 3. These results exhibit the superiority of our method and the importance of meticulously selecting the modalities for training in MFL.
Dataset CREMA-D AVE
setting IID non-IID IID non-IID
FedAvg 50.7 49.8 62.2 59.7
FedAvg-0.8 52.4 50.1 63.4 61.1
FedAvg-0.5 55.7 55.1 60.7 59.4
FedAvg-0.2 61.2 58.1 58.5 58.7
FedAvg+ML 55.8 54.5 62.8 60.7
DivFL+ML 57.1 55.6 63.0 61.1
CMSFed-local 63.7 60.3 63.4 60.1
CMSFed 64.5 61.6 64.7 62.1
Table 1: Ablation study. ‘PCE’ denotes using Eq. 9 for local enhancement. ‘CMSFed-local’ uses local prototypes rather than global prototypes.
CMSFed exploits all modalities comprehensively. To show the effect of our method on modality imbalance, we plot the accuracy curves of each modality on CREMA-D and AVE under IID setting. We can see from Fig. 5 that CMSFed could considerably improve the performance of weak modality (visual) and mitigate the modality-level bias as illustrated in Figs. 4(b) and 4(d). Besides, compared with randomly modal abandoning, which significantly reduces audio performance as illustrated in Fig. 1(b), CMSFed achieves comparable audio performance with other baselines. Since audio modality dominates multi-modal learning in these two datasets, pow-d and DivFL selects clients mainly relying on the loss or gradients from audio, which consequently improve audio performance but no improvement on visual modality. FedMSplit mainly focuses on modality incongruity, which is not an issue here, so there is no notable improvement in both modalities. Although FedProto, FedOGM and FedPMR also alleviate the imbalance, FedProto achieves worse audio behavior and FedOGM and FedPMR only focus on local optimization, resulting in the performance gap between them and our CMSFed on the aggregated model, which further indicates that in MFL, it is important to take both local optimization for each modality and the overall performance for aggregated model into consideration simultaneously.
5.4 Ablation study
Effectiveness of each component. Tab. 1 studies the effect of each CMSFed component. Applying Eq. 9 (FedAvg+PCE) surpasses the vanilla strategy (FedAvg) by a large margin, demonstrating its effectiveness on local enhancement. Comparing CMSFed (64.5% on IID CREMA-D) with ‘DivFL+PCE’ (57.1% on the same setting) also denotes the necessity of balancing different modalities via modality selection. To show the importance of aligning feature spaces of weak modality, we replace the global prototypes with local prototypes (CMSFed-local). Global alignment achieves notable improvement (by up to 2% on non-IID AVE). The performance improvement compared with ‘FedAvg-0.8,0.5,0.2’ exhibits that randomly sampling modalities does not always lead to improvement and further demonstrates the need of meticulously selecting modalities for information exploitation.
Figure 6: Proportional change of audio and visual respectively and the curve of global imbalance ratio during training on CREMA-D under IID setting.
The relationship between modality selection and imbalance degree. In Eqs. 13 and 14, we use a conflict resolution strategy based on local imbalance ratio to realize balanced modality selection. We visualize the proportions of audio and visual modalities selected in each round and the global imbalance ratio. It is clear from Fig. 6 that audio is the dominant modality (imbalance ratio is always greater than 1) and modality imbalance is gradually alleviated as training progresses (global imbalance ratio shows a downward trend). In addition, the proportion of selected visual modality follows the same trend (larger in the early stage of training and becomes smaller later), implying the rationality of our selection strategy based on imbalance ratio and its effectiveness on mitigating bias.
Effectiveness on modality incongruity scenario. All above experiments assume that each client initially has complete modal data. Here, we build the stimulation of modality incongruity scenario, in which half of the clients have data with two modalities, and the other half only have data in one modality: random audio or visual. The results are shown in Tab. 2 (The results of pow-d is not available because it is not applicable to this scenario). Our CMSFed still makes impressive improvement compared with all other baselines (by up to 3.5% on IID CREMA-D), illustrating the good generalization ability of our method in different scenarios. It is worth mentioning that FedMSplit performs better than before because it is specifically designed for modality incongruity.
Dataset CREMA-D AVE
setting IID non-IID IID non-IID
FedAvg 55.7 55.1 60.7 58.7
FedProx 56.8 56.0 61.2 58.5
FedProto 58.7 57.0 61.3 59.7
FedOGM 58.6 57.4 60.1 58.5
FedPMR 56.4 55.5 61.9 60.3
DivFL 57.4 55.6 61.1 58.6
FedMSplit 58.9 56.9 61.7 60.0
CMSFed 62.4 59.8 63.5 60.9
Table 2: Performance on CREMA-D and AVE with modality incongruity. 50% of clients have all modal data and 50% of clients only retain data with a single modality (audio or visual).
6 Conclusion
In this paper, we analyze traditional client selections in MFL and reveal there exists strong modality-level bias due to the modality imbalance during the training iterations. To address this issue, we propose the client-wise modality selection scheme for MFL (CMSFed) with modality-level gradient decoupling to release the potential of all modalities and maximize the gradient diversities to improve global aggregation. Our method does not introduce additional local training costs and communication overheads compared with previous methods. Extensive experiments on three datasets demonstrate the superiority of our method in performance and applicability under different modal combinations, data distributions and modality incongruity scenarios.

### v2

@@
-Client-wise Modality Selection for Balanced Multi-modal Federated Learning
+Overcome Modal Bias in Multi-modal Federated Learning via Balanced Modality Selection
 Abstract
-Selecting proper clients to participate in the iterative federated learning (FL) rounds is critical to effectively harness a broad range of distributed datasets.
-Existing client selection methods simply consider the variability among FL clients with uni-modal data, however, have yet to consider clients with multi-modalities.
-We reveal that traditional client selection scheme in MFL may suffer from a severe modality-level bias, which impedes the collaborative exploitation of multi-modal data, leading to insufficient local data exploration and global aggregation.
-To tackle this challenge, we propose a Client-wise Modality Selection scheme for MFL (CMSFed) that can comprehensively utilize information from each modality via avoiding such client selection bias caused by modality imbalance.
-Specifically, in each MFL round, the local data from different modalities are selectively employed to participate in local training and aggregation to mitigate potential modality imbalance of the global model.
-To approximate the fully aggregated model update in a balanced way, we introduce a novel local training loss function to enhance the weak modality and align the divergent feature spaces caused by inconsistent modality adoption strategies for different clients simultaneously.
-Then, a modality-level gradient decoupling method is designed to derive respective submodular functions to maintain the gradient diversity during the selection progress and balance MFL according to local modality imbalance in each iteration.
-Our extensive experiments showcase the superiority of CMSFed over baselines and its effectiveness in multi-modal data exploitation.
+Selecting proper clients to participate in each federated learning (FL) round is critical to effectively harness a broad range of distributed data.
+Existing client selection methods simply consider the mining of distributed uni-modal data, yet, their effectiveness may diminish in multi-modal FL (MFL) as the modality imbalance problem not only impedes the collaborative local training but also leads to a severe global modality-level bias.
+We empirically reveal that local training with a certain single modality may contribute more to the global model than training with all local modalities.
+To effectively exploit the distributed multiple modalities, we propose a novel Balanced Modality Selection framework for MFL (BMSFed) to overcome the modal bias.
+On the one hand, we introduce a modal enhancement loss during local training to alleviate local imbalance based on the aggregated global prototypes.
+On the other hand, we propose the modality selection aiming to select subsets of local modalities with great diversity and achieving global modal balance simultaneously.
+Our extensive experiments on audio-visual, colored-gray, and front-back datasets showcase the superiority of BMSFed over baselines and its effectiveness in multi-modal data exploitation.
+Keywords:
+Multi-modal federated learning modality imbalance modality selection
 1 Introduction
-Figure 1: Traditional client selection in FL aims to sample a client subset in each round.
-The selected clients use all their local data for training and aggregation.
-To exploit different modalities comprehensively, we propose the client-wise modality selection for MFL, which considers the local data from each modality as the sampling unit rather than the client.
-Only the selected modality data from a client would join in training and aggregation.
 Federated learning (FL) [23] aims to collaboratively learn data that has been collected by, and resides on, a number of remote devices or servers.
-FL stands to develop top-performing models by aggregating knowledge from numerous edge clients [6], which relies on the iterative interaction among participating clients and the server.
+FL stands to develop top-performing models by aggregating knowledge from numerous edge clients [8, 36], which relies on the iterative interaction among participating clients and the server.
 However, comprehensively employing the information from all clients can be exceptionally difficult due to the client heterogeneity and resource limitations [10, 29].
-Random sampling [32, 19] from available clients is widely used in FL to satisfy some practical restrictions, e.g., limited communication bandwidth [23, 31] and computing capacities [13].
-To improve the information exploitation of all clients, extensive research has been conducted on effective client selection strategies.
-One viable solution [24] is to gather as much local data as possible in each round while in contrast, some alternative studies [1] focus on collecting representative information by sampling a subset of clients.
-However, previous methods have only considered the knowledge discrepancies across different clients, while demonstrating notable performance in uni-modal FL, yet, their effectiveness diminishes when dealing with clients with multi-modal data as the inter-modality interactions during the FL training is neglected.
-The investigation into fine-grained and sophisticated data selection mechanism for multi-modal FL (MFL) is currently under-explored.
-In this paper, we reveal that traditional client selection approaches may suffer from a severe modality-level bias in MFL.
-Due to the inconsistent learning progress of different modalities, i.e., the modality imbalance, client-level selection bias tends to favor information from a dominant modality while deprioritize the weak modality, which impedes the collaborative exploitation of distributed multi-modal data, leading to insufficient local data exploration and global aggregation.
-To address such issue in MFL, we propose a novel Client-wise Modality Selection scheme for MFL (CMSFed), which chooses the appropriate modalities combination from clients in each round to alleviate the modality-level bias and realize full exploitation of each modality via considering the interactions between different modalities as illustrated in Fig.
+Random sampling [34, 22] from available clients has been widely used in FL to satisfy some practical restrictions, e.g., limited communication bandwidth [26, 35] and computing capacities [15].
+To improve the information exploitation of all clients, extensive research has been conducted on effective client selection strategies [9, 43].
+Despite the success of traditional client selection methods in uni-modal FL, their effectiveness diminishes when dealing with clients with multi-modal data as the inter-modal interactions during the MFL training are neglected.
+According to [37, 14], there may exist inconsistent learning paces for different modalities in multi-modal joint training, i.e., modality imbalance, which not only impedes the collaborative local training but also leads to a severe modal bias for global model in MFL.
+As illustrated in Tab.
+1, audio modality significantly outperforms visual modality in CREMA-D and AVE datasets during local training and the aggregated model still suffers from it.
+However, two well-designed client selection methods (pow-d [6] and DivFL [3]) only obtain severely limited improvement over random sampling (FedAvg) on the multi-modal global model.
+We can see that client selection methods achieve the best uni-audio performance while uni-visual performance even drops sometimes, which means existing client selection scheme heavily relies on the better modality, while ignoring the importance of improving weak modalities that also has potential for global model aggregation.
+Based on the above analysis, a pivotal question arises: Can we design a new selection scheme in MFL that can overcome the modal bias and exploit each modality comprehensively?
+Table 1: Performance of various client selection methods in MFL under IID setting.
+A and V denote uni-audio and uni-visual while A-V means the multi-modal result.
+‘Local’ represents that a client is trained based its local data without aggregation.
+A strong modal bias of global model exists on the two datasets.
+Dataset CREMA-D [4] AVE [33]
+Method A V A-V A V A-V
+Local 41.9 20.4 39.6 33.4 16.7 35.2
+FedAvg 51.2 20.6 50.7 61.1 26.8 62.2
+[…]
+DivFL [3] 52.3 21.1 51.7 62.7 25.3 63.3
+FedAvg-0.2 50.6 28.6 52.4 60.6 29.6 63.4
+FedAvg-0.5 50.5 34.6 55.7 58.7 30.0 60.7
+FedAvg-0.8 48.1 50.9 61.2 56.4 31.8 58.5
+BMSFed 51.0 41.9 64.5 59.7 40.2 64.7
+To answer this question, we investigate the interactions between different modalities via randomly discarding the data from one modality (audio or visual) on part of clients, which is inspired from modality dropout [2, 40] that drops a specific modal data during training for regularization.
+The results are shown in rows 7-9 in Tab.
+1, where ‘-x’ denotes randomly discarding a modality on a client with probability ‘x’.
+We can see that dropping with a certain probability can improve the global multi-modal performance on both datasets (e.g., FedAvg-0.2), and the main reason comes from the dramatic improvement of visual modality.
+This phenomenon suggests that performing uni-modal training can unleash its potential without being inhibited by another modality, and uni-modal local training may contribute more to the global model than multi-modal training on some clients .
+As the dropping ratio increases, although the visual modality still improves, multi-modal performance may decline as the audio modality declines as shown in columns 5 and 7, suggesting that we should carefully control which modalities on each client should be involved in training and aggregation to make the contribution most to the global model.
+Figure 1: Left: Traditional client selection in FL aims to sample a client subset in each round while our modality selection considers each local modality as the sampling unit.
+Right: The paradigm of BMSFed with four clients.
+The global prototypes are used to enhance the weak modality during local update.
+Only networks corresponding to the selected modalities will be uploaded to the server for aggregation.
+According to the above investigations, we propose a novel Balanced Modality Selection scheme for MFL (BMSFed) to mitigating the modal bias and comprehensively exploit the diverse information from all modalities.
+Specifically, instead of selecting a subset of clients, we treat each modality on the local side as a selection unit, as demonstrated in Fig.
 1.
-Only part of, not all, modalities, will be selected for some clients compared with traditional client selection.
-Specifically, in each communication round, we intend to gather a data subset by selectively employing different modalities from each client, the aggregated model update stemmed from that approximates the fully aggregated update which fulfills the potential of each modality.
-Inspired from [1] for diverse client selection, we build the modality selection formula, namely the min-form of the facility location function [5] evaluated on the collected data subset, to seek for the representative information for the global model.
-To further release the ability of each modality and reduce the training cost, we propose a modality-level gradient decoupling with a modal-balanced local training loss to build two separated submodular functions and then modulate their optimization process for unbiased modality selection.
-As a result, inconsistent modality adoption strategies may lead to different feature spaces for different clients, which exacerbates the difficulties in accomplishing an accurate and generalized aggregated model.
-Therefore, the global prototype is introduced for the weak modality to alleviate the heterogeneous feature spaces.
-To summarize, our contributions in this paper are as follows:
-  • We scrutinize the client selection scheme in MFL and reveal that naive sampling clients may introduce a strong modality-level bias which causes insufficient utilization of multi-modal data.
-  • We propose a novel client-wise modality selection scheme for MFL (CMSFed) to comprehensively exploit all modalities via a enhancing loss and gradient decoupling to balance MFL.
-  • We conduct comprehensive experiments and demonstrate that 1) CMSFed can achieve considerable improvements over baselines; 2) CMSFed indeed achieves a more balanced exploitation of distributed multimodal data.
+Our BMSFed mainly contains two parts: Firstly, we intend to alleviate the local modality imbalance by introducing a modal enhancement loss based on aggregated global prototypes to promote the performance of weak modality.
+Secondly, we complete the modality selection by building two separated submodular functions for selecting multi-modal clients (training with multi-modal data) and uni-modal clients (training with selected uni-modal data) respectively.
+Inspired from [3], the criterion is to select modalities that are most representative on the gradients while also alleviate the global modal bias.
+A simple yet effective conflict resolution strategy is devised to ensure the validity of modality selection and keep modal balance on global model simultaneously.
+The main contributions of the paper are summarized as follows:
+  1. 1.
+    We empirically analyze the modality imbalance problem in MFL and reveal that uni-modal training on some clients may contribute more to the global model than multi-modal training.
+  2. 2.
+    Based on the analysis, we propose a novel Balanced Modality Selection scheme for MFL (BMSFed) to comprehensively exploit all modalities via a modal enhancement loss and representative modality selection to overcome the global modal bias.
+  3. 3.
+    We conduct comprehensive experiments on audio-visual, colored-gray, and front-back datasets, and considering the statistical heterogeneity and modality incongruity problems in MFL, to validate the superiority of our BMSFed.
 2 Background and Related Works
-(a) Multi-modal
-(b) Uni-audio
-(c) Uni-visual
-Figure 2: Performance of client selection strategies on CREMA-D with uniformly distributed data on 20 clients.
-FedAvg randomly selects clients while DivFL selects clients with diversity.
-‘FedAvg-0.8’ means 80% of clients have data in both audio and visual modalities, while 20% of clients only have data in random one modality.
-‘-0.5’ and ‘-0.2’ follow the similar definition.
-(a) Performance of global multi-modal model.
-(b) Performance of global uni-audio model.
-(c) Performance of global uni-visual model.
 2.1 Multi-modal federated learning
-In MFL, each client may contain one or more modalities of data.
-Without loss of generality, we consider two input modalities, referred to as $A$ and $I$ in MFL.
-There are a set $V$ of $N$ clients, $V=\left[N\right]$, respectively owning datasets $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},y\right\}$ composed of $n_{i}$ pair samples, $i\in\left[N\right]$.
+In MFL, each client has one or multiple modalities of data.
+Without loss of generality, we consider two input modalities, which are denoted by $A$ and $I$ respectively in MFL.
+There are a set $V$ of $N$ clients, $V=\left[N\right]$, respectively owning datasets $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},\boldsymbol{y}\right\}$, $i\in\left[N\right]$.
 A typical federated learning objective is the average of each client’s local loss function:
-$\vspace{-8pt}\min f\left(\theta\right)=\sum_{k=1}^{N}{p_{k}F_{k}\left(\theta\right)}$ (1)
+$\vspace{-10pt}\min f\left(\theta\right)=\sum_{k=1}^{N}{p_{k}F_{k}\left(\theta\right)}$ (1)
 where $\theta=\left\{\theta^{A},\theta^{I},\omega\right\}$ denotes the model parameters.
-$\theta^{A}$ and $\theta^{I}$ represent the parameters of encoders of modality $A$ and $I$.
-$\omega$ is the parameters of fusion classifier.
+$\theta^{A}$ and $\theta^{I}$ represent the encoder parameters of modality $A$ and $I$.
+$\omega$ is the parameter of fusion classifier.
 $p_{k}$ is a pre-defined weight.
-$F_{k}$ is each client’s local loss.
+$F_{k}$ is each client’s local loss (cross entropy (CE) loss for classification task in this paper).
 Statistical heterogeneity [16, 20] is a widely concerned challenge in uni-modal FL.
@@
 The client selection via submodular maximization can be expressed following [3]: find a subset $S$ of clients whose aggregated gradients can approximate the full aggregation from all clients:
-$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}$ $=\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v^{\sigma\left(k\right)}\right)\right]}$ (2)
-$+\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}$
+$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}=\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v^{\sigma\left(k\right)}\right)\right]}+\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}$ (2)
 where $\sigma$ maps $V\rightarrow S$ and the gradient $\nabla F_{k}\left(v^{k}\right)$ from client $k$ is approximated by the gradient from a selected client $\sigma\left(k\right)\in S$.
@@
 Take the norms and apply triangular inequality after subtracting the second term from both sides, we can obtain a relaxed objective $G\left(S\right)$ for minimizing the approximation error:
-$\begin{split}&\left\|\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}-\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}\right\|\leqslant\\ &\sum_{k\in\left[N\right]}{\underset{i\in S}{\min}\left\|\nabla F_{k}\left(v^{k}\right)-\nabla F_{i}\left(v^{i}\right)\right\|\triangleq G\left(S\right)}\end{split}$ (3)
+$\begin{split}\left\|\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}-\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}\right\|\leqslant\sum_{k\in\left[N\right]}{\underset{i\in S}{\min}\left\|\nabla F_{k}\left(v^{k}\right)-\nabla F_{i}\left(v^{i}\right)\right\|\triangleq G\left(S\right)}\end{split}$ (3)
 Minimizing $G\left(S\right)$ can be seen as maximizing the well-known submodular function, i.e., the facility location function [7].
 The submodular maximizing problem is NP-hard but can be approximated via the greedy [25] or stochastic greedy algorithm [24]:
-$S\leftarrow S\cup k^{*},$ (4)
-$k^{*}\in$ $\underset{k\in\mathrm{rand}\left(V\backslash S,\mathrm{size}=s\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S\right)-\bar{G}\left(\left\{k\right\}\cup S\right)\right]$
+$S\leftarrow S\cup k^{*},k^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S,\mathrm{size}=s\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S\right)-\bar{G}\left(\left\{k\right\}\cup S\right)\right]$ (4)
 $\bar{G}$ represents a constant minus the negation of $G$.
-Although these methods make great improvement in uni-modal FL, the selection strategy is under-explored in MFL and we reveal that traditional client selection approaches fall into a severe bias in MFL.
+Although these methods make great improvement in uni-modal FL, the selection strategy is under-explored in MFL and we reveal that traditional client selection approaches cannot address the severe modal bias in MFL.
 2.3 Imbalanced multi-modal learning
@@
 In this paper, we aim to power each modality of all clients by a meticulously designed modality selection strategy in each round of training.
-3 Observations and Analysis
-In this section, we analyze the behavior of different client sampling strategies in MFL and the interactions between different modalities.
-As shown in Fig.
-2, we compare a well-designed client selection method (DivFL) with random sampling (FedAvg) on a multi-modal dataset.
-The results in Fig.
-1(a) show that the improvement from DivFL is severely limited although it is terrific in uni-modal FL.
-The results about how each modality behaves are illustrated in Figs.
-1(b) and 1(c).
-It is apparent that audio modality performs much better than visual modality and the optimised client selection options from DivFL generally focus on audio modality (audio performance of DivFL has significant improvement over FedAvg while no noticeable difference in visual).
-The performance gap between audio and visual can be attributed to ‘modality imbalance’ [33], which indicates the inconsistent learning paces in different modalities.
-Therefore, the imbalance nature (multi-modal learning is dominated by the modality who learns fast) [34] in MFL leads to a strong modality-level selection bias, i.e., client selection strategies tend to select clients that make the dominant modality aggregation better but neglect the weak modality
-To further analyze the interactions between different modalities, we randomly discard the data of one modality (audio or visual) on clients with certain probabilities (FedAvg-0.8, -0.5, -0.2) as demonstrated in Fig.
-1(a).
-It is clear that randomly discarding one modality data for part of clients does not necessarily lead to performance degradation of the global model, but may bring performance improvement.
-The reason behind can be found in Figs.
-1(b) and 1(c), where the gap between audio and visual modalities gradually narrows as the discarding probability increases until there is no stark disparity (in FedAvg-0.2).
-This phenomenon is consistent with the findings in [26]: the dominant modality hinders the learning process of weak modality.
-Therefore, the ability of visual modality is released and absorbed by the global model through aggregation when the audio data in some clients is abandoned.
-The above analyses suggest that traditional client selection methods are not applicable in MFL due to the modality imbalance nature, and the potential of each modality can be further explored by engaging appropriate modalities in training.
-Therefore, we intend to design a new modality selection scheme in MFL to address the modality imbalance and information underutilization from different modalities.
 3 Method
-In this section, we introduce CMSFed, a diverse and balanced modality selection strategy to realize more efficient data utilization.
-We first derive a combinatorial objective for modality selection in MFL via the gradient approximation with the fully aggregated update.
-Then we propose a modality-level gradient decoupling with the enhancing gradient on weak modality to build two separated submodular functions and balance MFL.
-4.1 Client-wise modality selection
-Firstly, we rewrite Eq. 2 for MFL as:
-$\begin{split}&\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)}=\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)}+\\ &\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v_{A}^{\sigma\left(k\right)},v_{I}^{\sigma\left(k\right)}\right)\right]}\end{split}$ (5)
-where $v_{A}^{k}$ and $v_{I}^{k}$ indicate the corresponding modal data in client $k$.
-The above equation is still the client selection formulation, while we could build the modality selection formula based on it as (data superscripts are omitted for simplicity):
-$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}$ (6)
-$=\sum_{k\in\left[N\right]}{\left[\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\nabla F_{\sigma_{M}\left(k\right)}\left(v_{A},v_{I}\right)\\ -\nabla F_{\sigma_{A}\left(k\right)}\left(v_{A}\right)-\nabla F_{\sigma_{I}\left(k\right)}\left(v_{I}\right)\\ \end{array}\right]}$
-$+\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}+\sum_{k\in S_{A}}{\gamma_{k}^{A}\nabla F_{k}\left(v_{A}\right)}$
-$+\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}$
-where $\sigma_{M}$, $\sigma_{A}$ and $\sigma_{I}$ map $V\rightarrow S_{M},S_{A},S_{I}$, the client sets who use multi-modal, $A$ modal, and $I$ modal data for training respectively, and $S_{M}\cap S_{A}=S_{A}\cap S_{I}=S_{M}\cap S_{I}=\oslash$.
-The full gradient aggregation is represented by the customized update from multi-modal and uni-modal clients.
-Assume modality $I$ is weak here, the relaxed objective for modality selection should be:
-$\left\|\begin{array}[]{c}\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}-\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}\\ -\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}\\ \end{array}\right\|$ (7)
-$\leqslant\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\\ \nabla F_{i}\left(v_{A},v_{I}\right)-\nabla F_{j}\left(v_{I}\right)\\ \end{array}\right\|}$
-uni-$A$ modal is omitted here as the multi-modal gradient is dominated by $A$, which means we do not need to select uni-$A$ clients for its acceleration.
-However, minimizing this objective, which is similar to client selection in Eq.
-3, cannot leverage sufficient information from each modality since the approximation objective (the left-hand side of Eq.
-6) still suffers from serious modality imbalance.
-How to finding a new full gradient aggregation that can exploit all modalities’ information is a great challenge.
-In addition to this, there also exist other challenges making Eq.
-7 difficult to be effectively optimized in practice:
-  • We need to build the relationship (a gradient similarity matrix) between fully aggregated gradient and local gradient for solving Eq.
-7.
-In uni-modal FL, a client only needs to train once to get gradients while in MFL, each client should perform local update twice (with multi-modal and weak-modal data respectively) to determine whether to choose multi-modal or uni-modal data for training.
-This inevitably leads to almost double the local training cost.
-  • Applying modality selection would cause different modality adoption strategies, namely modality heterogeneity, which results in divergent feature spaces across clients, especially on weak modality.
-To address the above challenges, we propose a modality-level gradient decoupling method with a balanced local loss, which not only leads to comprehensive information exploitation but also notably saves training overhead.
-4.2 Modality-level gradient decoupling
-We focus on the classification tasks in this paper, where cross-entropy (CE) loss is widely used. Hence, the loss for multi-modal and uni-modal learning should be:
-$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)$ (8)
-$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)$
-The naïve multi-modal loss falls into modality imbalance.
-Although obtaining a totally balanced solution seems impractical for multi-modal learning, we can find a surrogate for it: here we enhance the weak modality by equipping the naïve losses with the prototypical cross-entropy (PCE) loss $\beta^{k}\mathcal{L}_{P}^{k}$ [9] to correct gradient directions of weak modality:
-$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:\tilde{F}_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)$ (9)
-$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:\tilde{F}_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)+\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)$
-where $\mathcal{L}_{P}^{k}$ is calculated based on the local prototypes $c^{I}$:
-$\mathcal{L}_{P}^{k}\left(v_{I}\right)=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{I}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{I}\right)\right)}}\right]$ (10)
-where $d\left(\cdot,\cdot\right)$ is the distance function (Euclidean distance), $z_{i}^{I}$ is the feature of $x_{i}^{I}$.
+In this section, we introduce BMSFed that contains the local imbalance alleviation and balanced modality selection.
+3.1 Local Imbalance Alleviation
+As discussed in Sec.
+2.3 and Tab.
+1, the multi-modal training on each client may suffer from severe modality imbalance, leading to inadequate information exploitation and consequently incurring the modal bias on the global model.
+Therefore, we first try to alleviate the imbalance during local training.
+According to [11], the prototypes (i.e., class centroids) are suitable to calibrate the gradient directions to avoid the interference from other modalities.
+Therefore, we choose to use prototypes to facilitate learning of weak modality.
+For the $i$-th client with the data $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},\boldsymbol{y}\right\}$, the local prototype for class $j$ of each modality is defined as the mean value of representations:
+$c_{i,j}^{I}=\frac{1}{\left|\mathcal{D}_{i,j}\right|}\sum_{x^{I}\in\mathcal{D}_{i,j}}{h_{i}\left(\theta_{i}^{I};x^{I}\right)},c_{i,j}^{A}=\frac{1}{\left|\mathcal{D}_{i,j}\right|}\sum_{x^{A}\in\mathcal{D}_{i,j}}{h_{i}\left(\theta_{i}^{A};x^{A}\right)}$ (5)
+where $\mathcal{D}_{i,j}$ denotes the samples belonging to $j$-th class in client $i$.
+$h_{i}$ is the function of encoder.
+Considering the heterogeneity across clients, we aggregate the local prototypes to a global prototype as:
+$c_{j}^{GI}=\frac{1}{\left|\mathcal{N}_{j}\right|}\sum_{i\in\mathcal{N}_{j}}{\frac{\left|\mathcal{D}_{i,j}\right|}{N_{j}}c_{i,j}^{I}},c_{j}^{GA}=\frac{1}{\left|\mathcal{N}_{j}\right|}\sum_{i\in\mathcal{N}_{j}}{\frac{\left|\mathcal{D}_{i,j}\right|}{N_{j}}c_{i,j}^{A}}$ (6)
+where $\mathcal{N}_{j}$ denotes the set of clients that have class $j$ and $N_{j}$ is the number of instances belonging to class $j$ over all clients.
+Then, we introduce a modal enhancement loss (ME) based on global prototype to adjust local training:
+$\mathcal{L}_{ME}^{k}\left(v_{I}^{k}\right)$ $=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{GI}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{GI}\right)\right)}}\right]$ (7)
+$\mathcal{L}_{ME}^{k}\left(v_{A}^{k}\right)$ $=-\mathbb{E}_{\left(x_{i}^{A},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{A},c_{y}^{GA}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{A},c_{j}^{GA}\right)\right)}}\right]$
+where $d\left(\cdot,\cdot\right)$ is the distance function (Euclidean distance), $z_{i}^{I}$ is the representation of $x_{i}^{I}$, i.e., $z_{i}^{I}=h_{i}\left(\theta_{i}^{I};x_{i}^{I}\right)$.
 $Y$ is the class number.
-The calculation details for $c^{I}$, $\beta^{k}$ are in Appendix.
-Applying $\mathcal{L}_{P}^{k}$ on uni-modal loss is also essential, which can improve this modality by clustering features tightly.
-Bring Eq. 9 to Eq. 7 and we can get the new objective:
-$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\tilde{F}_{k}\left(v_{A},v_{I}\right)-\\ \nabla\tilde{F}_{i}\left(v_{A},v_{I}\right)-\nabla\tilde{F}_{j}\left(v_{I}\right)\\ \end{array}\right\|}=$ (11)
-$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\nabla\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{i}\left(v_{A},v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{P}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{P}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
+$v_{A}^{k}$ and $v_{I}^{k}$ indicate the corresponding modal data in client $k$.
+Hence, the local loss should be (data superscripts are omitted for simplicity):
+$F_{k}\left(v_{A},v_{I}\right)=\left\{\begin{array}[]{c}\begin{matrix}\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\gamma^{k}\mathcal{L}_{ME}^{k}\left(v_{A}\right),&\rho_{I}^{k}\leqslant 1\\ \end{matrix}\\ \begin{matrix}\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right),&\rho_{I}^{k}>1\\ \end{matrix}\\ \end{array}\right.$ (8)
+where $\rho_{I}^{k}$ indicates the local imbalance ratio for client $k$ calculated based on the modality-wise ground truth prediction, and $\rho_{I}^{k}>1$ means modality $A$ outperforms modality $I$ and vice verse.
+$\gamma^{k},\beta^{k}\in\left(0,1\right)$ are the modulation coefficients.
+The calculation details for $\gamma^{k}$, $\beta^{k}$ and $\rho_{I}^{k}$ are in Appendix.
+Different from [11], we apply the global prototype instead of local prototypes to aggregate the knowledge from different clients, and the ME loss is not only applied to multi-modal clients, but also to uni-modal clients (See details in Sec.
+3.2).
+Now, the weak modality will be stimulated through local update.
+3.2 Balanced Modality Selection
+Considering that different types of modal combinations may be selected to participate in local training in our scheme, we define the participated clients as multi-modal clients and uni-modal clients (e.g., client 2 in Fig.
+1 Right is a uni-modal client with modal $I$ for training while client 1 and 4 are multi-modal clients with both modalities for training).
+In Sec.
+3.1, we propose the ME loss to facilitate the local learning of weak modality, which is originally designed for multi-modal clients.
+Here, we also apply it on uni-weak-modal clients to realize gradient consistency on different types of clients.
+Assume modality $I$ is weak here. Hence, the local loss for multi-modal and uni-modal clients should be:
+$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)$ (9)
+$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{A}\right)=\mathcal{L}_{CE}^{k}\left(v_{A}\right),F_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)+\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)$
+Next, following Eq.
+2, we define the paradigm of modality selection for MFL, aiming to approximate the full gradient aggregation from all clients by the gradient from the customized multi-modal and uni-modal clients:
+$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}=\sum_{k\in\left[N\right]}{\left[\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\nabla F_{\sigma_{M}\left(k\right)}\left(v_{A},v_{I}\right)\\ -\nabla F_{\sigma_{A}\left(k\right)}\left(v_{A}\right)-\nabla F_{\sigma_{I}\left(k\right)}\left(v_{I}\right)\\ \end{array}\right]}$ (10)
+$+\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}+\sum_{k\in S_{A}}{\gamma_{k}^{A}\nabla F_{k}\left(v_{A}\right)}+\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}$
+where $\sigma_{M}$, $\sigma_{A}$ and $\sigma_{I}$ map $V\rightarrow S_{M},S_{A},S_{I}$, the client sets who use multi-modal, uni-$A$, and uni-$I$ data for training respectively, and $S_{M}\cap S_{A}=S_{A}\cap S_{I}=S_{M}\cap S_{I}=\oslash$.
+$\gamma_{k}^{M}$, $\gamma_{k}^{A}$ and $\gamma_{k}^{I}$ have the similar meaning as $\gamma_{k}$ in Eq.
+2.
+Since modality $I$ is weak here, we omit the uni-$A$ clients as the multi-modal gradient is dominated by modality $A$ [28], which means we do not need to select uni-$A$ clients for its enhancement.
+Then, bring Eq.
+9 to Eq.
+10 and follow the operations from Eq.
+2 to Eq.
+3, we can obtain:
+$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\nabla F_{k}\left(v_{A},v_{I}\right)-{\gamma_{i}^{M}}\nabla F_{i}\left(v_{A},v_{I}\right)-{\gamma_{j}^{I}}\nabla F_{j}\left(v_{I}\right)\right\|}$ (11)
+$=\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\nabla\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)-\nabla\mathcal{L}_{CE}^{i}\left(v_{A},v_{I}\right)\\ -\nabla\beta^{i}\mathcal{L}_{ME}^{i}\left(v_{I}\right)-\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{ME}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
 $\leqslant\sum_{k\in\left[N\right]}{\underset{i\in S_{M}}{\min}\left\|\nabla\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)-\nabla\mathcal{L}_{CE}^{i}\left(v_{A},v_{I}\right)\right\|}$
-$+\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{P}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{P}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
+$+\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{ME}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{ME}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
 $\triangleq G\left(S_{M}\right)+G\left(S_{M}\cup S_{I}\right)$
-Firstly, replacing Eq.
-8 with Eq.
-9 can enhance the weak modality on the local side.
-However, the right-hand side of the first equation is still a complex issue.
-Then, we decouple this objective into two submodular functions according to triangle inequality and the full gradient approximation is divided into two parts: the first part uses selected multi-modal CE gradient to fit fully multi-modal CE gradient aggregation and the second part approximates the fully multi-modal PCE gradient aggregation via selected uni-modal CE gradient and both selected multi- and uni-modal PCE gradient.
-Two gradient similarity matrices are required for two parts.
-The modality-level gradient decoupling not only reduces algorithm complexity (convert joint selection to two selection problems) but also highly lowers training cost as each selected client only needs to perform once local training according to its selected modality type to update gradient similarity matrices (both matrices for multi-modal clients, only the second matrix for uni-modal clients).
-After modality selection, the gradient of $\theta^{I}$ from each client varies dramatically since there exists an obvious difference between the multi-modal learning gradient and uni-modal learning gradient for the weak modality, as discussed in [9].
-To address this heterogeneity, we replace the local prototype in Eq.
-10 with the global prototype $c^{GI}$:
-$\mathcal{L}_{GP}^{k}\left(v_{I}\right)=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{GI}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{GI}\right)\right)}}\right]$ (12)
-Applying the global PCE loss to Eq. 9 can enhance local weak modality and align the feature spaces of various clients simultaneously.
-Then, we can perform modality selection with the stochastic greedy algorithm for two submodular functions:
-$S_{M}\leftarrow S_{M}\cup k_{1}^{*},$ (13)
-$k_{1}^{*}\in$ $\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S_{M}\right)-\bar{G}\left(\left\{k\right\}\cup S_{M}\right)\right]$
-Since the second submodular function involves both $S_{M}$ and $S_{I}$, we add a simple yet effective conflict resolution strategy to ensure $S_{M}\cap S_{I}=\oslash$ as well as, more importantly, balance the learning of different modalities:
-$\begin{cases}if\,\,k_{1}^{*}=k_{2}^{*},S_{M}\cup k_{2}^{*};\\ if\,\,k_{1}^{*}\neq k_{2}^{*},\left\{\begin{array}[]{c}S_{I}\cup k_{2}^{*},if\,\,\rho_{I}^{k}>\chi\\ S_{M}\cup k_{2}^{*},if\,\,\rho_{I}^{k}\leqslant\chi\\ \end{array}\right.\\ \end{cases}$ (14)
-$k_{2}^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\begin{array}[]{c}\bar{G}\left(S_{M}\cup S_{I}\right)-\\ \bar{G}\left(\left\{k\right\}\cup S_{M}\cup S_{I}\right)\\ \end{array}\right]$
-where $\rho_{I}^{k}$ is the imbalanced ratio of client $k$ as in [9], which is the quotient of average ground-truth logits from two modalities (See the Appendix for details).
-$\rho_{I}^{k}>1$ means audio is dominant while visual is dominant when $\rho_{I}^{k}<1$.
-$\chi$ is a hyper-parameter for threshold.
-It tends to select weak modality for severely imbalanced clients to alleviate imbalance, while selected balanced clients use multi-modal data to improve overall performance.
+The right-hand side of the first equation is the modality selection formula that aims to select a group of multi-modal clients $S_{M}$ and uni-modal clients $S_{I}$ to approximate the aggregated gradients from all clients.
+However, the joint selection for $S_{M}$ and $S_{I}$ is a complex joint optimization problem.
+Therefore, we decouple this objective into two submodular functions $G\left(S_{M}\right)$ and $G\left(S_{M}\cup S_{I}\right)$ according to triangle inequality, while the full gradient approximation is divided into two parts: the first part uses selected multi-modal CE gradient to fit fully multi-modal CE gradient aggregation, and the second part approximates the fully multi-modal ME gradient aggregation via selected uni-modal CE gradient and both selected multi- and uni-modal ME gradient.
+The modality-level gradient decoupling converts the complex joint selection to two simply separated selection problems.
 Input: Input data $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},y\right\}$, $i\in\left[N\right]$, initial model $\theta$, hyper-parameters $\chi$, global communication epochs $E$, $e=1$.
@@
       else
-            Aggregate global model, prototypes and $\rho_{I}$;
-      Select a subset of modalities $S_{M}$ and $S_{I}$ (or $S_{A}$) using Eqs. 13 and 14;
-       Send $\theta$ to selected clients;
+            Aggregate global model $\theta$, prototypes $c^{G}$ and $\rho_{I}$;
+      Select multi-modality for $S_{M}$ and uni-modality for $S_{I}$ (or $S_{A}$) using Eqs. 12 and 13;
+       Send $\theta$, $c^{G}$ and $\rho_{I}$ to selected clients;
        foreach client in selected clients in parallel do
-             Perform multi-modal learning in $S_{M}$ and uni-modal learning in $S_{I}$ (or $S_{A}$) by Eq. 9 with Eq. 12;
-             Send gradients, prototypes and $\rho_{I}^{k}$ to server;
-Algorithm 1 CMSFed.
-(a) CREMA-D
-(b) AVE
-(c) CG-MNIST
-Figure 3: Test accuracy of CMSFed compared with other baselines on CREMA-D, AVE and CG-MNIST under IID settings. CMSFed converges to more accurate solutions than all baselines in these three datasets.
-(a) CREMA-D
-(b) AVE
-(c) CG-MNIST
-Figure 4: Test accuracy of CMSFed compared with other baselines on CREMA-D, AVE and CG-MNIST under non-IID settings. CMSFed still makes remarkable improvement.
+             Perform multi-modal learning in $S_{M}$ and uni-modal learning in $S_{I}$ (or $S_{A}$) by with Eqs. 8 and 9;
+             Send gradients, local prototypes and $\rho_{I}^{k}$ to server;
+Algorithm 1 BMSFed.
+Although we can solve the two submodular functions with the stochastic greedy algorithm [24], there are still two issues: (1) the selected client according to $G\left(S_{M}\cup S_{I}\right)$ should be specified whether it is uni-modal client or multi-modal client; (2) the separated selection strategy pays less attention to the global modal bias.
+To address the two problems, we perform the stochastic greedy algorithm for two submodular functions in parallel and propose a simple yet effective conflict resolution strategy to ensure $S_{M}\cap S_{I}=\oslash$ as well as, more importantly, balance the learning of different modalities on global model:
+$S_{M}\leftarrow S_{M}\cup k_{1}^{*},k_{1}^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S_{M}\right)-\bar{G}\left(\left\{k\right\}\cup S_{M}\right)\right]$ (12)
+$\begin{cases}if\,\,k_{1}^{*}=k_{2}^{*},S_{M}\cup k_{2}^{*};\\ if\,\,k_{1}^{*}\neq k_{2}^{*},\left\{\begin{array}[]{c}S_{I}\cup k_{2}^{*},if\,\,\rho_{I}^{k}>\chi\\ S_{M}\cup k_{2}^{*},if\,\,\rho_{I}^{k}\leqslant\chi\\ \end{array}\right.\\ \end{cases}$ (13)
+$k_{2}^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S_{M}\cup S_{I}\right)-\bar{G}\left(\left\{k\right\}\cup S_{M}\cup S_{I}\right)\right]$
+For every selection, we randomly sample a subset of clients $s$.
+A multi-modal clients $k_{1}^{*}$ is selected from $s$ according to $G\left(S_{M}\right)$ while $k_{2}^{*}$ is also selected from $s$ according to $G\left(S_{M}\cup S_{I}\right)$.
+$k_{1}^{*}=k_{2}^{*}$ means using multi-modal data from this client can contribute most to the global model.
+When $k_{1}^{*}\neq k_{2}^{*}$, we allocate it to uni-modal or multi-modal client according to its local imbalance ratio: if it is severely imbalanced, we use its uni-weak-modal data for training and aggregation to alleviate the global modal bias, otherwise we believe that training with its multi-modal data contributes more than uni-modal data.
+$\chi$ is a hyper-parameter.
 Discussion.
-(1) Mitigating the selection bias in our method are twofold: the PCE loss $\tilde{F}_{k}$ alleviates imbalance at local side and the selected uni-modal clients further promote balanced learning of global model.
-Meanwhile, the diversity coming from two submodular functions ensures the improvement for the global model.
-(2) The maximization of $\bar{G}\left(S_{M}\right)$ and $\bar{G}\left(S_{M}\cup S_{I}\right)$ can be viewed as the optimization for multi-modal learning and uni-weak-modal learning respectively.
-Imbalance ratio is adopted to keep their balance.
-(3) We assume $I$ is the weak modality above while in practice, we can determine the weak modality before modality selection via the aggregated global imbalance ratio $\rho_{I}$ with $\rho_{I}^{k}$ from all clients.
-Overall, the pseudo-code of CMSFed is provided in Algorithm 1.
+(1) Overcoming the modal bias in our method are twofold: the ME loss alleviates imbalance at local side and the selected uni-modal clients further promote balanced learning of global model.
+Meanwhile, the diversity coming from two submodular functions ensures the representative information for the global model.
+(2) We assume $I$ is the weak modality above while in practice, we can determine the weak modality before modality selection via the aggregated global imbalance ratio $\rho_{I}=\frac{1}{\sum\nolimits_{k=1}^{N}{n_{k}}}\sum\nolimits_{k=1}^{N}{\rho_{I}^{k}\cdot n_{k}}$.
+Overall, the pseudo-code of BMSFed is provided in Algorithm 1.
 (3) Only the gradients, prototypes and $\rho_{I}^{k}$ participate in communication, so there is no privacy issue and similar communication overheads as in traditional FL.
@@
 4.1 Datasets and baselines
-CREMA-D [2] is an audio-visual dataset for emotion recognition task, each video in which consists of both facial and acoustic emotional expressions.
-There are total 6 categories for emotional states: neutral, happy, sad, fear, disgust and anger.
-7,442 clips in total are collected in this dataset.
-6,698 samples are randomly chosen as the training set and the rest of 744 samples are the testing set.
-AVE [30] is an audio-visual dataset for audio-visual event localization, which consists of 28 event classes and 4,143 10-second video clips with both auditory and visual tracks as well as second-level annotations.
-All the video clips are collected from YouTube.
-In our experiments, we aim to construct a labeled multi-modal classification dataset by extracting the frames from event-localized video segments and capturing the audio clips within the same segments.
-The training and validation splits of the dataset follow [30].
-Colored-and-gray MNIST [15] is a synthetic dataset based on MNIST [16], and we denote it as CG-MNIST in this paper.
-Two kinds of images form a sample pair: a gray-scale image and a monochromatic image.
-In the training set, there are 60,000 sample pairs, and the monochromatic images are strongly color-correlated with their digit labels, In the validation set, the number of sample pairs is 10,000, while the monochromatic images are weakly color-correlated with their labels.
-The data synthesis method follows [34].
-This dataset is used to prove the method’s effectiveness beyond audio-visual modality.
-(a) Audio in CREMA-D
-(b) Visual in CREMA-D
-(c) Audio in AVE
-(d) Visual in AVE
-Figure 5: The performance of each modality compared with other baselines on CREMA-D and AVE under IID settings.
-CMSFed achieves comparable audio accuracy and significantly better visual performance simultaneously.
+Datasets.
+We conduct experiments on four datasets: (1) CREMA-D [4] is an audio-visual dataset for emotion recognition task with total six categories for emotional states.
+(2) AVE [33] is an audio-visual dataset for event localization with 28 event classes, and here we use it to construct a labeled multi-modal classification dataset following [11].
+(3) Colored-and-gray MNIST (CG-MNIST) [17] is a synthetic dataset based on MNIST [18] with gray-scale and monochromatic images as two modalities, following [38].
+(4) ModelNet40 is one of the Princeton ModelNet datasets [39] with 3D objects of 40 categories.
+The front and back [31] views are considered as two modalities, following [5].
+Details of these four datasets are in the Appendix.
 Baselines.
@@
 4.2 Experimental settings
-For CREMA-D and AVE, we use ResNet18 [10] as the backbone for both audio and visual modalities.
+For CREMA-D, AVE and ModelNet40, we use ResNet18 [12] as the backbone for audio, visual and flow modalities.
 Audio data is converted to a spectrogram of size 257x299 for CREMA-D and 257×1,004 for AVE.
@@
 For CG-MNIST, we build a neural network with 4 convolution layers and 1 average pool layer as the encoder, following the setting as in [11].
-We choose the simple yet effective fusion method, concatenation [25], to build fusion classifier for the three datasets.
-We set 20 clients for CREMA-D, AVE and 30 clients for CG-MNIST, while 5 clients are selected in each communication round for CREMA-D, AVE and 6 for CG-MNIST.
-For IID setting, training data is uniformly distributed to clients.
-For non-IID scenarios, we use Dirichlet distribution [11] $Dir\left(\alpha\right)$ to split data ($\alpha=3$ for CREMA-D and AVE, $\alpha=2$ for CG-MNIST).
+We choose the simple yet effective fusion method, concatenation [27], to build fusion classifier for all the datasets.
+We set 20 clients for CREMA-D, AVE and ModelNet40 while the number for CG-MNIST is 30.
+5 clients are selected in each communication round for CREMA-D, AVE, ModelNet40 and 6 for CG-MNIST.
+For IID setting, training data is uniformly distributed to all clients.
+For non-IID scenarios, we use Dirichlet distribution [13] $Dir\left(\alpha\right)$ to split data ($\alpha=3$ for CREMA-D, AVE, ModelNet40, $\alpha=2$ for CG-MNIST).
 The optimizer is SGD [30] for all datasets.
-Learning rate is initialized at 1e-3 or 1e-2 for CEAMA-D and AVE or CG-MNIST and becomes 1e-4 or 1e-3 in the later training stage.
-The hyper-parameter $\chi$ is set to 1.5-2.5 according to datasets and settings.
+Learning rate is initialized at 1e-3 or 1e-2 for CEAMA-D, AVE and ModelNet40 or CG-MNIST and becomes 1e-4 or 1e-3 in the later training stage.
+The hyper-parameter $\chi$ is set to 1.2-2.5 according to datasets and settings.
 To complete stochastic greedy algorithm for Eqs.
-13 and 14, we use the gradients from the selected clients at current round to update part of the similarity matrix, which is named “no-overheads” in [1].
-Except for pow-d, DivFL and CMSFed, other baselines select clients randomly.
-Each client has two modal data by default if not specified.
+12 and 13, we use the gradients from the selected clients at current round to update part of the similarity matrix, which is named “no-overheads” in [3].
+Except for pow-d, DivFL and BMSFed, other baselines select clients randomly.
+Each client has two modal data by default.
+We do all experiments on a workstation with an RTX 3090 GPU, a 3.9-GHZ Intel Core i9-12900K CPU and 64GB of RAM.
+Table 2: Comparison results on four datasets.
+The metric is the top-1 accuracy (%).
+The best is in bold, and the second best is underlined.
+Our method achieves significant improvement on both IID and non-IID scenarios.
+Dataset CREMA-D AVE CG-MNIST ModelNet40
+Method IID non-IID IID non-IID IID non-IID IID non-IID
+FedAvg 50.7 49.8 62.2 59.7 42.3 41.7 87.2 86.5
+FedProx 51.0 49.0 62.6 59.9 42.9 43.6 86.9 87.1
+FedProto 58.7 54.0 61.7 58.8 51.5 51.4 87.5 87.2
+FedOGM 56.9 56.4 62.8 59.3 57.2 53.0 87.6 87.0
+FedPMR 55.5 55.1 63.1 61.6 66.1 63.3 87.6 87.7
+pow-d 50.5 50.7 62.5 60.0 41.2 40.3 86.8 86.2
+DivFL 51.7 50.8 63.3 59.6 43.0 42.1 86.5 86.4
+FedMSplit 52.4 51.6 62.4 60.8 43.5 50.9 87.5 87.4
+BMSFed 64.5 61.6 64.7 62.1 70.2 66.7 88.7 87.5
+(a) CREMA-D under IID
+(b) AVE under IID
+(c) CREMA-D under non-IID
+(d) AVE under non-IID
+Figure 2: Test accuracy of BMSFed compared with other baselines on CREMA-D and AVE. BMSFed converges to more accurate solutions than all baselines.
+Table 3: The uni-modal performance comparison on CREMA-D and AVE.
+The metric is the top-1 accuracy (%).
+Our method achieves the best uni-visual (the weak modality) performance on all settings and maintains the comparable uni-audio performance.
+Dataset CREMA-D AVE
+Setting IID non-IID IID non-IID
+Method A V A V A V A V
+FedAvg 51.2 20.6 50.7 20.2 61.1 26.8 61.4 26.4
+FedProx 51.3 20.2 50.1 22.0 60.4 27.1 61.2 26.9
+FedProto 50.2 35.3 48.6 39.1 55.7 36.8 59.7 32.8
+FedOGM 50.5 35.7 48.8 30.2 58.7 28.8 59.4 29.4
+FedPMR 51.5 38.7 50.1 35.9 61.7 39.6 61.7 35.3
+pow-d 51.5 20.4 51.6 18.8 61.9 26.9 60.1 27.1
+DivFL 52.3 21.1 52.1 22.7 62.7 25.3 61.6 26.3
+FedMSplit 52.0 21.8 50.8 21.6 61.3 26.9 62.3 28.7
+BMSFed 51.0 41.9 49.3 41.4 59.7 40.2 60.2 38.6
 4.3 Comparison with baselines
-CMSFed effectively improves the performance.
-As shown in Fig.
-3, we report the test accuracy versus the number of communication rounds for three datasets under IID data distributions.
-We observe that CMSFed can achieve significant improvement compared with all other baselines (by up to 5% on CREMA-D).
-CMSFed also realizes comparable or even faster convergence speeds in CREMA-D and CG-MNIST.
-The slow learning at the beginning in AVE is attributed to the need for more effort to explore the information of weak modality.
+BMSFed effectively improves the performance.
+As demonstrated in Tab.
+2, our BMSFed achieves the best results on the four datasets under both IID and non-IID settings (by up to 5.8% on CREMA-D).
 Client sampling here (pow-d and DivFL) cannot fully exploit information for all modalities, making its improvement limited or even worse than FedAvg in CG-MNIST.
-Although FedOGM and FedPMR accomplish modest improvement because of their ability to alleviate modality imbalance, they are not as good as CMSFed since they do not consider the overall performance of aggregated model.
+Although FedOGM and FedPMR accomplish modest improvement because of their ability to alleviate modality imbalance, they are not as good as BMSFed since they do not consider the overall performance of the global model.
 Traditional uni-modal FL methods for statistical heterogeneity (e.g.
 FedProx) and MFL method for modality incongruity (FedMSplit) only obtain slight improvement.
-The results for non-IID data distributions are demonstrated in Fig.
-4.
-The overall trends are similar to the results in Fig.
+We also illustrate the trend of test accuracy versus the number of communication rounds on CREMA-D and AVE in Fig.
+2.
+BMSFed can realize comparable or even faster convergence speeds in CREMA-D and AVE.
+BMSFed exploits all modalities comprehensively.
+To show the effect of our method on addressing modal bias, we report the performance of each modality on CREMA-D and AVE as shown in Tab.
 3.
-These results exhibit the superiority of our method and the importance of meticulously selecting the modalities for training in MFL.
+The uni-modal performance evaluation follows [11]: a sample is classified into the class corresponding to its nearest prototype.
+It is clear that BMSFed could considerably improve the performance of weak modality (visual) and mitigate the modality-level bias.
+Besides, compared with randomly modality abandoning, which significantly reduces audio performance as illustrated in Tab.
+1, BMSFed achieves comparable audio performance with other baselines.
+Although FedProto, FedOGM and FedPMR also alleviate the imbalance, they mainly focus on local optimization, resulting in the performance gap between them and our BMSFed on the aggregated model, which further indicates that in MFL, it is important to take both local optimization for each modality and the overall performance for global model into consideration simultaneously.
+Table 4: Ablation study. ‘BMSFed-local’ uses local prototypes rather than global prototypes for ME loss.
 Dataset CREMA-D AVE
@@
 FedAvg 50.7 49.8 62.2 59.7
-FedAvg-0.8 52.4 50.1 63.4 61.1
+DivFL 51.7 50.8 63.3 59.6
+FedAvg-0.2 52.4 50.1 63.4 61.1
 FedAvg-0.5 55.7 55.1 60.7 59.4
-FedAvg-0.2 61.2 58.1 58.5 58.7
-FedAvg+ML 55.8 54.5 62.8 60.7
-DivFL+ML 57.1 55.6 63.0 61.1
-CMSFed-local 63.7 60.3 63.4 60.1
-CMSFed 64.5 61.6 64.7 62.1
-Table 1: Ablation study. ‘PCE’ denotes using Eq. 9 for local enhancement. ‘CMSFed-local’ uses local prototypes rather than global prototypes.
-CMSFed exploits all modalities comprehensively.
-To show the effect of our method on modality imbalance, we plot the accuracy curves of each modality on CREMA-D and AVE under IID setting.
-We can see from Fig.
-5 that CMSFed could considerably improve the performance of weak modality (visual) and mitigate the modality-level bias as illustrated in Figs.
-4(b) and 4(d).
-Besides, compared with randomly modal abandoning, which significantly reduces audio performance as illustrated in Fig.
-1(b), CMSFed achieves comparable audio performance with other baselines.
-Since audio modality dominates multi-modal learning in these two datasets, pow-d and DivFL selects clients mainly relying on the loss or gradients from audio, which consequently improve audio performance but no improvement on visual modality.
-FedMSplit mainly focuses on modality incongruity, which is not an issue here, so there is no notable improvement in both modalities.
-Although FedProto, FedOGM and FedPMR also alleviate the imbalance, FedProto achieves worse audio behavior and FedOGM and FedPMR only focus on local optimization, resulting in the performance gap between them and our CMSFed on the aggregated model, which further indicates that in MFL, it is important to take both local optimization for each modality and the overall performance for aggregated model into consideration simultaneously.
+FedAvg-0.8 61.2 58.1 58.5 58.7
+FedAvg+ME 55.8 54.5 62.8 60.7
+DivFL+ME 57.1 55.6 63.0 61.1
+BMSFed-local 63.7 60.3 63.4 60.1
+BMSFed 64.5 61.6 64.7 62.1
 4.4 Ablation study
@@
 Tab.
-1 studies the effect of each CMSFed component.
-Applying Eq.
-9 (FedAvg+PCE) surpasses the vanilla strategy (FedAvg) by a large margin, demonstrating its effectiveness on local enhancement.
-Comparing CMSFed (64.5% on IID CREMA-D) with ‘DivFL+PCE’ (57.1% on the same setting) also denotes the necessity of balancing different modalities via modality selection.
-To show the importance of aligning feature spaces of weak modality, we replace the global prototypes with local prototypes (CMSFed-local).
+4 studies the effect of each BMSFed component.
+Applying ME loss Eq.
+7 on random sampling (FedAvg+ME) and the well-designed client selection (DivFL+ME) surpasses their vanilla strategies (FedAvg, DivFL) by a large margin, demonstrating its effectiveness on local enhancement.
+Comparing BMSFed (64.5% on IID CREMA-D) with ‘DivFL+ME’ (57.1% on the same setting) also denotes the necessity of balancing different modalities considering the global model via modality selection.
+To show the importance of aligning feature spaces of weak modality, we replace the global prototypes with local prototypes (BMSFed-local).
 Global alignment achieves notable improvement (by up to 2% on non-IID AVE).
-The performance improvement compared with ‘FedAvg-0.8,0.5,0.2’ exhibits that randomly sampling modalities does not always lead to improvement and further demonstrates the need of meticulously selecting modalities for information exploitation.
+The performance improvement compared with ‘FedAvg-0.2,0.5,0.8’ exhibits that randomly sampling modalities does not always lead to improvement and further demonstrates the need of meticulously selecting modalities for information exploitation.
+(a) CREMA-D under IID
+(b) CREMA-D under IID
+(c) CG-MNIST under IID
+Figure 3: Robustness validation on data size, local epoch and client number. Our BMSFed consistently outperforms baseline (FedAvg) under various scenarios.
 Figure 4: Proportional change of audio and visual respectively and the curve of global imbalance ratio during training on CREMA-D under IID setting.
+Robustness test.
+To verify the robustness of our method, we vary three key hyperparameters to build various scenarios: (1) change the data size $\left|\mathcal{D}_{i}\right|$ to allow each client to hold a small amount of data, (2) set different local training epochs, and (3) vary the total client number $N$.
+As illustrated in Figs.
+3(a) and 3(b), our BMSFed consistently outperforms baseline (FedAvg) under various scenarios.
+More data as well as more local training epochs can bring further improvements to our method, indicating that exploiting the weak modality need more training efforts.
+Based on the results in Fig.
+3(c), our approach can also be generalised to clients with larger scales.
 The relationship between modality selection and imbalance degree.
 In Eqs.
-13 and 14, we use a conflict resolution strategy based on local imbalance ratio to realize balanced modality selection.
+12 and 13, we use a conflict resolution strategy based on local imbalance ratio to realize balanced modality selection.
 We visualize the proportions of audio and visual modalities selected in each round and the global imbalance ratio.
@@
 5 (The results of pow-d is not available because it is not applicable to this scenario).
-Our CMSFed still makes impressive improvement compared with all other baselines (by up to 3.5% on IID CREMA-D), illustrating the good generalization ability of our method in different scenarios.
+Our BMSFed still makes impressive improvement compared with all other baselines (by up to 3.5% on IID CREMA-D), illustrating the good generalization ability of our method in different scenarios.
 It is worth mentioning that FedMSplit performs better than before because it is specifically designed for modality incongruity.
+Table 5: Performance on CREMA-D and AVE with modality incongruity. 50% of clients have all modal data and 50% of clients only retain data with a single modality (audio or visual).
 Dataset CREMA-D AVE
@@
 FedMSplit 58.9 56.9 61.7 60.0
-CMSFed 62.4 59.8 63.5 60.9
-Table 2: Performance on CREMA-D and AVE with modality incongruity. 50% of clients have all modal data and 50% of clients only retain data with a single modality (audio or visual).
+BMSFed 62.4 59.8 63.5 60.9
 5 Conclusion
-In this paper, we analyze traditional client selections in MFL and reveal there exists strong modality-level bias due to the modality imbalance during the training iterations.
-To address this issue, we propose the client-wise modality selection scheme for MFL (CMSFed) with modality-level gradient decoupling to release the potential of all modalities and maximize the gradient diversities to improve global aggregation.
+In this paper, we analyze traditional client selections and find their ineffectiveness in MFL.
+We further reveal that there exists strong modality-level bias due to the modality imbalance during the training iterations and uni-modal training on some clients may contribute more to the global model than multi-modal training.
+To address this issue, we propose the balanced modality selection scheme for MFL (BMSFed) with modality-level gradient decoupling to release the potential of all modalities and maximize the gradient diversities to improve global aggregation.
+We also introduce a modal enhancement loss to optimize the local update process.
 Our method does not introduce additional local training costs and communication overheads compared with previous methods.
-Extensive experiments on three datasets demonstrate the superiority of our method in performance and applicability under different modal combinations, data distributions and modality incongruity scenarios.
+Extensive experiments on four datasets demonstrate the superiority of our method in performance and applicability under different modal combinations, data distributions and modality incongruity scenarios.
+Acknowledgements
+Our work was supported by two grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No.
+PolyU15222621, PolyU15225023) and the funding from National Natural Science Foundation of China under grants 62302184.